6. At 60°C, pure substance A has a vapor pressure of 380 mmHg and pure substance B of 140 mmHg. The mole fraction XB=0.15 is in the range of validity of Henry's law. The partial pressure of B at XB=0.1 is 20 mmHg. (a) What is the vapor partial pressure PB at XB=0.15? Answer: PB=30 mmHg. (b) What would PB be at XB=0.15 if the solution were ideal? Answer: PB=21 mmHg. (c) What is the activity coefficient of B, γB, at XB=0.15, using pure liquid B as the standard state for B? Answer: 1.428. (d) What is the molar free energy change ∆GM at XB=0.15 after mixing A and B? Answer: ∆GM= -1022.78j. (e) Is the attraction between A and B stronger or weaker than the attraction between A and A, and B and B? Explain your answer.
Answer #1

a) Given PAo = 380 mm Hg ; PBo = 140 mm Hg
By using Dalton' s law , Ptotal * XB = PB
   Ptotal * 0.1 = 20 mm Hg
   Ptotal = 200 mm Hg
at XB = 0.15 ; PB = 200 * 0.15 = 30 mm Hg

b) If the solution were ideal, just use the Raoult's law to solve
   PB = XB * PBo
   PB = 0.15 * 140 mm Hg = 21 mm Hg

c) at XB = 0.15 ;
   γB = PB / PBo = 200 / 140 = 1.428571

d) molar mixing free energy is = ΔGm = RT (XA ln XA + XB ln XB )
XA = 1 - 0.15 = 0.85
  
ΔGm = 8.314 * (60+273.15) [ 0.85 * ln(0.85) + 0.15* ln(0.15) ] = -1170.823 J

e) as , γB > 1 the interactions between B-B repel each other, as the pure vapor pressure is increases in A
   the interactions between A-A will be stronger than B-B. Pure V.P exerted by A and B will be more than
Individual A and B itself so, the attractive forces between A-B will be strongest as its activity coeffcient
   decreases , which is usually less than zero.
NOTE:
If the interactions attract each other then γ<1, If they repel one another then γ>1
If there are no interactions then γ=1